Use the information below to answer the following 2 questions.

A rubber-band slingshot is drawn back is drawn back 0.15 m with a force of 28 N and shoots a 28-g stone. Solve using energy considerations. Recall from the previous lesson that Ep = ½ kx2. Since F = kx, Ep = ½ Fx.


1.The initial speed of the stone if the rubber band is released is___m/s (Give your answer with the correct number of sig digs and do not include units).


2.The maximum height of the stone if it is shot straight up is ___m. (Give your answer with the correct number of sig digs and do not include units).


1
Expert's answer
2021-07-19T09:49:17-0400

1)


"0.5Fx=0.5mv^2\\(28)(0.15)=(0.028)v^2\\v=12\frac{m}{s}"


2)


"0.5Fx=mgh\\0.5(28)(0.15)=(0.028)(9.8)h\\h=7.7\ m"